Gaseous nitrogen actuates a Carnot power cycle in which the respective volumes at the four corners of the cycle, starting at the beginning of the isothermal expansion, are 𝑉1 = 10.10 𝐿, 𝑉2 = 14.53 𝐿, 𝑉3 = 226.54 𝐿, 𝑉4 = 157.73 𝐿. The cycle receives 21.1 KJ of heat. Determine the work.


1
Expert's answer
2021-11-25T11:51:27-0500

"V_1 = 10.1 \;L \\

V_2 = 14.53 \;L \\

V_3 = 226.54 \;L \\

V_4 = 157.73 \;L"

It is the case of reverce carnot cycle.

Heat resieved by cycle is

"Q_L = 21.1 \;kJ"

For process 2-3:

Heat interaction is

"Q_L = -W

21.1 \times 10^3 = nRT_2ln(\frac{V_3}{V_2}) \\

21.1 \times 10^3 = 1 \times \frac{8314}{28} \times T_2 \times ln(\frac{226.54}{14.53}) \\

T_2 = 25.87 \;K"

Now, for process 3-4 is an adiabatic process

"\frac{T_H}{T_L} = (\frac{V_3}{V_4})^{Y-1} \\

Y=1.4"

for adiabatic

"\frac{T_H}{25.87} = (\frac{226.54}{157.73})^{0.4} \\

T_H = 29.901 \;K"

Now, heat transfer in process 1-4

"Q_{14} = -nRTln(\frac{V_1}{V_4}) \\

= -1 \times \frac{8314}{28} \times 29.901 \times ln(\frac{10.1}{157.73}) \\

= 24.401 \; kJ"

Net heat transfer "= Q_{12} +Q_{23} +Q_{34} +Q_{41}"

"= 0-21.1+0+24.401 \\

= 3.301 \;kJ"

In a cycle we know that

"\sum Q_{net} = \sum W_{net} \\

W_{net} = 3.301 \;kJ"